OpenStudy (anonymous):

solve x^3/2=4 1 fan 1 vote ;)

OpenStudy (anonymous):

I already answered this for you, but I can explain it again :)

OpenStudy (anonymous):

You have x^3/2 = 4 Your first step is to get rid of the division This can be done by multiplying each side by 2 This will give you: x^3 = 8 Are you with me so far?

OpenStudy (anonymous):

Now you have to cube root each side By cube rooting x^3, you get rid of the eponant x = cuberoot(8) x = 2 (2x2x2) = 8 2^3 = 8

OpenStudy (anonymous):

thx mate :D